# Constraint on initial conditions of one-dimensional expanding fluids from nonlinear causality

### v1

@@ Abstract
The initial conditions of one-dimensional expanding viscous fluids in relativistic heavy-ion collisions are scrutinized in terms of nonlinear causality of the relativistic hydrodynamic equations. Conventionally, it is believed that the matter generated in relativistic heavy-ion collisions starts to behave as a fluid all at once at some initial time. However, it is by no means trivial how soon after the first contact of two high-energy nuclei the fluid picture can be applied. It is demonstrated that one-dimensional expanding viscous fluids violate the necessary and the sufficient conditions of nonlinear causality at large departures from local equilibrium. We therefore quantify the inverse Reynolds number to justify the hydrodynamic description to be valid. The initial conditions are strictly constrained not to violate the causality conditions during the time evolution. With the help of the transverse energies per rapidity measured at RHIC and LHC, we obtain the minimum initial proper time and the maximum energy density allowed by nonlinear causality. This analysis strongly suggests that the initial stage of relativistic heavy-ion collisions needs to be described by a non-equilibrium framework other than relativistic dissipative hydrodynamics.
@@ I Introduction
A vast body of experimental data at the Large Hadron Collider (LHC) in CERN and at Relativistic Heavy-Ion Collider (RHIC) in Brookhaven National Laboratory (BNL) together with theoretical analysis of these data by using sophisticated dynamical models have revealed rich and profound physics of the quark gluon plasma (QGP). The QGP is a deconfined nuclear matter which occupied the early Universe $\sim 10$ micro seconds after the Big Bang and its temperature reaches $\sim 10^{12}$ Kelvin. Hence the physics of the QGP elucidates a state of the matter composed of fundamental degree of freedom at extremely high-temperature Yagi et al. (2005).
To extract bulk and transport properties of the QGP, relativistic hydrodynamics has been playing a vital role as a framework to describe dynamics of local equilibrium states. Since the QGP is created as a transient state of matter in relativistic heavy-ion collisions, it is of crucial importance to develop an integrated dynamical model based on relativistic hydrodynamics which describes multi-stages, in particular, before and after hydrodynamic evolution. In the final state of high-energy nuclear collisions, the description of dynamics can be changed from macroscopic relativistic hydrodynamics to microscopic kinetic theory for dilute hadron gases. This became a standard prescription in dynamical modeling once the importance of the hadronic corona was recognized Hirano and Gyulassy (2006). On the other hand, it has not been established so far how to describe the pre-hydrodynamic stage, namely how to initialize hydrodynamic fields in relativistic hydrodynamics. The hydrodynamic equations as first order differential equations with respect to time demand initial conditions. Conventionally, one either parametrizes the initial hydrodynamic fields so that final momentum distributions provide a reasonable reproduction of experimental data or employs a framework to describe the early stages of dynamics until the hydrodynamic initial time. It is known that the initial conditions of hydrodynamic models cannot be uniquely determined. One of the most reliable ways is to constrain the initial conditions through Bayesian parameter estimation Bernhard et al. (2016, 2019); Auvinen et al. (2020); Nijs et al. (2021a, b); Parkkila et al. (2021); Everett et al. (2021); Heffernan et al. (2023); Shen et al. (2023). However, a lack of the knowledge to describe the dynamics prior to hydrodynamic evolution could bring a systematic uncertainty of posterior distributions of the parameters even by using Bayesian estimation.
@@
The second-order hydrodynamics including Muller–Israel–Stewart type constitutive equations Müller (1967); Israel (1976); Israel and Stewart (1979) are known to obey causality as long as the relaxation time is sufficiently large. This was shown for the first time in Ref. Hiscock and Lindblom (1983) within propagation of a linear perturbation around a static and uniform local equilibrium background in which energy density and pressure are constant ($e=$ const. and $p=$ const.), flow velocity vanishes ($u^{\mu}=(1,0,0,0)$), and the dissipative currents such as shear stress tensor and bulk pressure vanish ($\pi^{\mu\nu}=0$ and $\Pi=0$).
However, the essential aspects of the second-order terms cannot be captured in the linear perturbation since these dissipative terms are higher order in small quantities and neglected in the linear regime. Since the second-order hydrodynamic equations are a set of quasi-linear partial differential equation, causal propagation and local stability can be discussed without relying on small deviation. Therefore it is of particular importance to go beyond the linear regime and to capture full nonlinearity of relativistic dissipative hydrodynamic equations. To our best knowledge, the first study of nonlinear causality in the second-order relativistic hydrodynamic equations was done in one-dimensional expanding system in Ref. Hiscock and Olson (1989). Later some attempts were made to investigate a relation between causality and stability of the solutions of nonlinear hydrodynamic equations Denicol et al. (2008); Pu et al. (2010); Floerchinger and Grossi (2018). Recently, necessary and sufficient conditions are obtained from full nonlinear hydrodynamic equations in the nonlinear regime Bemfica et al. (2021). Since these conditions are written as inequalities including thermodynamic variables and dissipative currents, these can be used to (in-)validate the numerical solutions of hydrodynamics equations Plumberg et al. (2022); Chiu and Shen (2021). These could provide a new theoretical guidance on relativistic hydrodynamic simulations and robust Bayesian inference towards quantitative extraction of the properties of the QGP da Silva et al. (2023); Krupczak et al. (2024); Domingues et al. (2024).
After a brief summary of the second-order hydrodynamics in Sec. II, we show the necessary and the sufficient conditions, constraint on initial conditions, and the equation-of-state dependence of one-dimensional expanding fluids in Sec. III. Section IV is devoted to summary and conclusion of the present study. We also discuss more details about the one-dimensional expanding conformal fluids in Appendix A and review the necessary and the sufficient conditions of nonlinear causality in Appendix B.
@@ II Model
Relativistic hydrodynamics conventionally describes space-time evolution of thermodynamic variables, dissipative currents, and four fluid velocity under an assumption that the local system is close to thermal equilibrium. One can also regard its framework as an effective theory to describe long-time and long-range behaviors for a given microscopic system based on the conservation laws of energy and momentum and the gradient expansion. In the current framework of the second order relativistic hydrodynamic, dissipative currents are promoted to dynamical variables. Therefore one has to scrutinize whether given initial conditions, in particular, for dissipative currents are within a valid range of the framework or not.
@@ II.1 Relativistic Hydrodynamics
where $T^{\mu\nu}$ is the energy-momentum tensor, $e$ is the energy density, $p$ is the pressure, $u^{\mu}$ is the four fluid velocity, and $\pi^{\mu\nu}$ is the shear stress tensor. The four fluid velocity is normalized as $u_{\mu}u^{\mu}=1$. By decomposing $T^{\mu\nu}$ into the time-like and the space-like components for each Lorentz index, these can be defined as
@@
In principle, one might choose any initial conditions for a differential equation of the first order in time. On the other hand, the shear stress tensor $\pi^{\mu\nu}$ is one of the dissipative currents so as to be small in comparison with thermodynamic pressure. Thus it is by no means trivial whether any initial conditions for $\pi^{\mu\nu}$ can be chosen or not. This is exactly the issue which we are going to address in the next section from a viewpoint of nonlinear causality.
@@ II.2 Equation of State
In this study, we consider two models of equation of state (EoS): the equation of state of conformal system (conformal EoS) and the equation of state from numerical calculations of lattice QCD (lattice EoS). We employ these two models for the sake of comparison to see effects of nonconformality on final results.
For the conformal system, the trace of the energy-momentum tensor vanish, $T^{\mu}_{\enskip\mu}=e-3p=0$. Therefore the conformal EoS is [^1] Since this EoS is also obtained by assuming the ideal gas system of relativistic particles, it is sometimes recognized that interactions among constituent particles are neglected in this particular EoS. However, it is not necessary to be true. Neglecting interactions among the constituent particles is not necessary to obtain this EoS. Only an assumption that particles are massless leads to the conformal system and, consequently, $p=e/3$.
@@
where $\bar{t}=T/(0.154\enskip\rm{GeV})$ is the dimensionless temperature, $p_{\rm{id}}=95\pi^{2}/180$ is a prefactor of pressure for an ideal gas at the Stefan–Boltzmann limit including the degree of freedom for three flavor QCD, $c_{t}=3.8706$, $t_{0}=-0.9761$, $a_{n}=-8.7704$, $b_{n}=3.9200$, $d_{n}=0.3419$, $a_{d}=-1.26$, $b_{d}=0.8425$, and $d_{d}=-0.0475$ are the fitting parameters Bazavov et al. (2014). The other thermodynamic quantities and the square of sound velocity ($c_{s}^{2}=dp/de$) are obtained from Eq. (8) through the thermodynamic relations. In this way, the lattice EoS, $p=p(e)$, is obtained.
@@ II.3 Equation of Motion in One-Dimensional Expanding System
In what follows, we assume translational and rotational symmetries in the transverse plane perpendicular to the collision axis and boost invariance in the longitudinal direction. Then we employ the boost invariant solution of four flow velocity Bjorken (1983), $u^{\mu}_{\mathrm{Bj}}=(t,0,0,z)/\tau$, where $\tau=\sqrt{t^{2}-z^{2}}$ is the proper time. From Eqs. (1) and (4), one obtains Baier et al. (2008)
@@
where $\phi\equiv\pi^{00}-\pi^{33}$ is the “shear pressure”. From a translational and rotational symmetries in the transverse plane, the shear stress tensor at local rest frame in one-dimensional boost invariant system is
@@ II.4 Time Evolution of the Inverse Reynolds Number
Figure 1 shows the time evolution of the dimensionless nonequilibrium quantity, $\phi/(e+p)$, from the solution of the hydrodynamic equations (9) and (10) for (a) the conformal EoS and (b) the lattice EoS. Since the nonlinear term ($\propto\phi^{2}$) in Eq. (10) was neglected in several studies (see, e.g., Ref. Heller and Spalinski (2015)), the solution with this nonlinear term (solid lines) is compared with the one without it (dashed lines). Initial times are increased incrementally from $\tau/\tau_{\pi}=0.5$ to $4.5$ at intervals of 1.0 for $\phi/(e+p)=\pm 0.6$. In the case of lattice QCD EoS, one needs to fix the energy scale since it is no longer conformal. Therefore, as initial conditions, we choose $e=5$ GeV/fm3, calculate the pressure $p=p(e=5\mbox{GeV/fm}^{3})$ and choose the shear pressure so that it becomes $\phi/(e+p)=\pm 0.6$.
@@
It is well-known that the solutions of Eqs. (9) and (10) rapidly converge to a single curve which is sometimes called “hydrodynamic attractor” Heller and Spalinski (2015); Romatschke (2017, 2018); Florkowski et al. (2018); Jankowski and Spaliński (2023). It is obvious from Fig. 1 that all results exhibit this behavior, regardless of whether one includes the nonlinear term in Eq. (10). In fact, we also exhibit some solutions leave away from hydrodynamic attractor if we start from too small $\phi/(e+p)$ in Appendix A.
@@ III Results
We first obtain nonlinear causal conditions in one-dimensional expanding system within the BRSSS equation in Sec. III.1. For given transport coefficients from AdS/CFT, the necessary and the sufficient conditions can be translated into the constraints on the dimensionless nonequilibrium quantity (or the inverse Reynolds number) in Sec. III.2. Even if the system starts from the local equilibrium state, the system can be deviated from it due to strong longitudinal expansion and may have a large value of the inverse Reynolds number. This means the initial conditions can be chosen so that the system obeys the causality even during its evolution. This will be done in Sec. III.3. In Sec. III.4, we discuss how the results with the conformal EoS are modified when we employ the lattice EoS.
@@ III.1 Nonlinear Causality in One-Dimensional Expanding System
The sufficient conditions of one-dimensional expanding fluids from the BRSSS equation (4) depend on the sign of the shear pressure $\phi$. Regardless of the sign of $\phi$, almost trivial inequality of the EoS and the transport coefficients are obtained from the sufficient conditions:
@@ III.2 Constraint on the Inverse Reynolds Number
By employing conformal EoS and transport coefficients from AdS/CFT, the causality conditions are written as inequalities among thermodynamic variables and the shear pressure: [^3] Although the pressure $p$ is just equal to $e/3$ in the conformal EoS, we still keep $p$ in these inequalities since we compare these results with the one from lattice EoS in Sec. III.4.  [^4] Out result is consistent with the one obtained in Ref. Bemfica et al. (2021): $-4/5<\Lambda_{a}/(e+p)\leq 3/5$ with a different choice of relaxation time, $\tau_{\pi}T=5\eta/s$.
@@
As for the necessary conditions, Eqs. (18) and (19) provide the constraint (31) for $\phi<0$ and $\phi>0$, respectively. On the other hand, the constraint (32) comes from Eqs. (III.1) and (III.1). See also Appendix B to understand how each necessary condition is obtained from a discussion of characteristic velocity.
Figure 2: Examples of the solutions of hydrodynamic equations (9) and (10) for several initial conditions and acausal (red) and causal (blue) regions from the necessary (31) and the sufficient (32) conditions, respectively. Note that one cannot determine whether it is causal or not in between only from these conditions.
To see how the solutions for hydrodynamic equations of one-dimensional expanding fluids obey or violate the causality, the regions from these conditions (31) and (32) together with ten examples of solutions are shown in Fig. 2. When a solution of Eqs. (9) and (10) goes through the acausal region, it does not obey at least one necessary condition and, consequently, violates the relativistic causality. On the other hand, when a solution keeps its trajectory in the causal region during the time evolution, it obeys the all sufficient conditions (needless to say, the all necessary conditions, too) and, consequently, the causality of the dynamics is guaranteed. Only from these necessary and sufficient conditions of nonlinear causality, it is not yet determined in between whether it is causal or not: All necessary conditions are satisfied, but some of the sufficient conditions are not satisfied. In any cases, the dynamics of fluids tends to be acausal when the inverse Reynolds number is large and the system is far away from local equilibrium.

### v2

@@ Abstract
The initial conditions of one-dimensional expanding viscous fluids in relativistic heavy-ion collisions are scrutinized in terms of nonlinear causality of the relativistic hydrodynamic equations. Conventionally, it is believed that the matter generated in relativistic heavy-ion collisions starts to behave as a fluid all at once at some initial time. However, it is by no means trivial how soon after the first contact of two high-energy nuclei the fluid picture can be applied. It is demonstrated that one-dimensional expanding viscous fluids violate the necessary and the sufficient conditions of nonlinear causality at large departures from local equilibrium. We therefore quantify the inverse Reynolds number to justify the validity of the hydrodynamic description. The initial conditions are strictly constrained not to violate the causality conditions during the time evolution. With the help of the transverse energies per rapidity measured at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC), we obtain the minimum initial proper time and the maximum energy density allowed by nonlinear causality. This analysis strongly suggests that the initial stage of relativistic heavy-ion collisions needs to be described by a non-equilibrium framework other than relativistic dissipative hydrodynamics.
@@ I Introduction
A vast body of experimental data at the Large Hadron Collider (LHC) in CERN and at Relativistic Heavy-Ion Collider (RHIC) in Brookhaven National Laboratory (BNL) together with theoretical analysis of these data by using sophisticated dynamical models have revealed rich and profound physics of the quark gluon plasma (QGP). The QGP is a deconfined nuclear matter which occupied the early Universe $\approx 10$ micro seconds after the big bang and its temperature reaches $\approx 10^{12}$ Kelvin. Hence the physics of the QGP elucidates a state of the matter composed of fundamental degree of freedom at extremely high temperature Yagi et al. (2005).
To extract bulk and transport properties of the QGP, relativistic hydrodynamics has been playing a vital role as a framework to describe dynamics of local equilibrium states. Since the QGP is created as a transient state of matter in relativistic heavy-ion collisions, it is of crucial importance to develop an integrated dynamical model based on relativistic hydrodynamics which describes multiple stages, in particular, before and after hydrodynamic evolution. In the final state of high-energy nuclear collisions, the description of dynamics can be changed from macroscopic relativistic hydrodynamics to microscopic kinetic theory for dilute hadron gases. This became a standard prescription in dynamical modeling once the importance of the hadronic corona was recognized Hirano and Gyulassy (2006). On the other hand, it has not been established so far how to describe the pre-hydrodynamic stage, namely how to initialize hydrodynamic fields in relativistic hydrodynamics. The hydrodynamic equations as first order differential equations with respect to time demand initial conditions. Conventionally, one either parametrizes the initial hydrodynamic fields so that final momentum distributions provide a reasonable reproduction of experimental data or employs a framework to describe the early stages of dynamics until the hydrodynamic initial time. It is known that the initial conditions of hydrodynamic models cannot be uniquely determined. One of the most reliable ways is to constrain the initial conditions through Bayesian parameter estimation Bernhard et al. (2016, 2019); Auvinen et al. (2020); Nijs et al. (2021a, b); Parkkila et al. (2021); Everett et al. (2021); Heffernan et al. (2023); Shen et al. (2023). However, a lack of the knowledge to describe the dynamics prior to hydrodynamic evolution could bring a systematic uncertainty of posterior distributions of the parameters even by using Bayesian estimation.
@@
The second-order hydrodynamics including Muller–Israel–Stewart type constitutive equations Müller (1967); Israel (1976); Israel and Stewart (1979) are known to obey causality at least within a linear regime as long as the relaxation time is sufficiently large. This was shown for the first time in Ref. Hiscock and Lindblom (1983) within propagation of a linear perturbation around a static and uniform local equilibrium background in which energy density and pressure are constant ($e=$ const. and $p=$ const.), flow velocity vanishes [$u^{\mu}=(1,0,0,0)$], and the dissipative currents such as shear stress tensor and bulk pressure vanish ($\pi^{\mu\nu}=0$ and $\Pi=0$).
However, the essential aspects of the second-order terms cannot be captured in the linear perturbation since these dissipative terms are higher order in small quantities and neglected in the linear regime. Since the second-order hydrodynamic equations are a set of quasi-linear partial differential equation, causal propagation and local stability can be discussed without relying on small deviation. Therefore it is of particular importance to go beyond the linear regime and to capture full nonlinearity of relativistic dissipative hydrodynamic equations. To our best knowledge, the first study of nonlinear causality in the second-order relativistic hydrodynamic equations was done in one-dimensional expanding system in Ref. Hiscock and Olson (1989). Later some attempts were made to investigate a relation between causality and stability of the solutions of nonlinear hydrodynamic equations Denicol et al. (2008); Pu et al. (2010); Floerchinger and Grossi (2018). Recently, necessary and sufficient conditions were obtained from full nonlinear hydrodynamic equations in the nonlinear regime Bemfica et al. (2021). Since these conditions are written as inequalities including thermodynamic variables and dissipative currents, these can be used to (in-)validate the numerical solutions of hydrodynamics equations Plumberg et al. (2022); Chiu and Shen (2021). These could provide new theoretical guidance on relativistic hydrodynamic simulations and robust Bayesian inference towards quantitative extraction of the properties of the QGP da Silva et al. (2023); Krupczak et al. (2024); Domingues et al. (2024).
After a brief summary of the second-order hydrodynamics in Sec. II, we show the necessary and the sufficient conditions, constraints on initial conditions, and the equation-of-state dependence of one-dimensional expanding fluids in Sec. III. Section IV is devoted to a summary and conclusion of the present study. We also discuss more details about the one-dimensional expanding conformal fluids in Appendix A and review the necessary and the sufficient conditions of nonlinear causality in Appendix B.
@@ II Model
Relativistic hydrodynamics conventionally describes space-time evolution of thermodynamic variables, dissipative currents, and fluid four-velocity under an assumption that the local system is close to thermal equilibrium. One can also regard its framework as an effective theory to describe long-time and long-range behaviors for a given microscopic system based on the conservation laws of energy and momentum and the gradient expansion. In the current framework of the second-order relativistic hydrodynamics, dissipative currents are promoted to dynamical variables. Therefore one has to scrutinize whether given initial conditions, in particular, for dissipative currents are within a valid range of the framework or not.
@@ II.1 Relativistic hydrodynamics
where $T^{\mu\nu}$ is the energy-momentum tensor, $e$ is the energy density, $p$ is the pressure, $u^{\mu}$ is the fluid four-velocity, and $\pi^{\mu\nu}$ is the shear stress tensor. The fluid four-velocity is normalized as $u_{\mu}u^{\mu}=1$. By decomposing $T^{\mu\nu}$ into the time-like and the space-like components for each Lorentz index, these can be defined as
@@
In principle, one might choose any initial conditions for a differential equation of the first order in time. On the other hand, the shear stress tensor $\pi^{\mu\nu}$ is one of the dissipative currents, in order to be small in comparison with thermodynamic pressure. Thus it is by no means trivial whether any initial conditions for $\pi^{\mu\nu}$ can be chosen or not. This is exactly the issue that we are going to address in the next section from a viewpoint of nonlinear causality.
@@ II.2 Equation of state
In this study, we consider two models of the equation of state (EoS): the equation of state of a conformal system (conformal EoS) and the equation of state from numerical calculations of lattice QCD (lattice EoS). We employ these two models for the sake of comparison to see effects of nonconformality on final results.
For the conformal system, the trace of the energy-momentum tensor vanishes, $T^{\mu}_{\enskip\mu}=e-3p=0$. Therefore the conformal EoS is [^1] Since this EoS is also obtained by assuming the ideal gas system of relativistic particles, it is sometimes recognized that interactions among constituent particles are neglected in this particular EoS. However, it is not necessary for this assumption to be true. Neglecting interactions among the constituent particles is not necessary to obtain this EoS. Only an assumption that particles are massless leads to the conformal system and, consequently, $p=e/3$.
@@
where $\bar{t}=T/(0.154\enskip\rm{GeV})$ is the dimensionless temperature, $p_{\rm{id}}=95\pi^{2}/180$ is a prefactor of pressure for an ideal gas at the Stefan–Boltzmann limit including the degree of freedom for three flavor QCD, and $c_{t}=3.8706$, $t_{0}=-0.9761$, $a_{n}=-8.7704$, $b_{n}=3.9200$, $d_{n}=0.3419$, $a_{d}=-1.26$, $b_{d}=0.8425$, and $d_{d}=-0.0475$ are the fitting parameters Bazavov et al. (2014). Although these parameters are obtained by fitting the lattice QCD results in the range $0.40.13<T<$ 0.4 GeV Bazavov et al. (2014), we utilize them in the whole temperature region. The other thermodynamic quantities and the square of sound velocity ($c_{s}^{2}=dp/de$) are obtained from Eq. (8) through the thermodynamic relations. In this way, the lattice EoS, $p=p(e)$, is obtained.
@@ II.3 Equation of Motion in One-Dimensional Expanding System
In what follows, we assume translational and rotational symmetries in the transverse plane perpendicular to the collision axis and boost invariance in the longitudinal direction. Then we employ the boost invariant solution of flow four-velocity Bjorken (1983), $u^{\mu}_{\mathrm{Bj}}=(t,0,0,z)/\tau$, where $\tau=\sqrt{t^{2}-z^{2}}$ is the proper time. From Eqs. (1) and (4), one obtains Baier et al. (2008)
@@
where $\phi\equiv\pi^{00}-\pi^{33}$ is the “shear pressure”. From the translational and rotational symmetries in the transverse plane, the shear stress tensor for the local rest frame in the one-dimensional boost invariant system is
@@ II.4 Time Evolution of the Inverse Reynolds Number
Figure 1 shows the time evolution of the dimensionless nonequilibrium quantity, $\phi/(e+p)$, from the solution of the hydrodynamic equations (9) and (10) for (a) the conformal EoS and (b) the lattice EoS. Since the nonlinear term ($\propto\phi^{2}$) in Eq. (10) was neglected in several studies (see, e.g., Ref. Heller and Spalinski (2015)), the solution with this nonlinear term (solid lines) is compared with the one without it (dashed lines). Initial times are increased incrementally from $\tau/\tau_{\pi}=0.5$ to $4.5$ at intervals of 1.0 for $\phi/(e+p)=\pm 0.6$. In the case of the lattice QCD EoS, one needs to fix the energy scale since it is no longer conformal. Therefore, as initial conditions, we choose $e=5$ GeV/fm3, calculate the pressure $p=p(e=5$ GeV/fm3) and choose the shear pressure so that it becomes $\phi/(e+p)=\pm 0.6$.
@@
It is well-known that the solutions of Eqs. (9) and (10) rapidly converge to a single curve which is sometimes called “hydrodynamic attractor” Heller and Spalinski (2015); Romatschke (2017, 2018); Florkowski et al. (2018); Jankowski and Spaliński (2023). It is obvious from Fig. 1 that all results exhibit this behavior, regardless of whether one includes the nonlinear term in Eq. (10). In fact, in AppendxA we also exhibit some solutions that move away from the hydrodynamic attractor if we start from too small $\phi/(e+p)$.
@@ III Results
We first obtain nonlinear causal conditions in one-dimensional expanding system within the BRSSS equation in Sec. III.1. For given transport coefficients from AdS/CFT, the necessary and sufficient conditions can be translated into the constraints on the dimensionless nonequilibrium quantity (or the inverse Reynolds number) in Sec. III.2. Even if the system starts from the local equilibrium state, the system can deviated from it due to strong longitudinal expansion and may have a large value of the inverse Reynolds number. This means the initial conditions can be chosen so that the system obeys the causality even during its evolution. This will be done in Sec. III.3. In Sec. III.4, we discuss how the results with the conformal EoS are modified when we employ the lattice EoS.
@@ III.1 Nonlinear Causality in One-Dimensional Expanding System
The sufficient conditions of one-dimensional expanding fluids from the BRSSS equation (4) depend on the sign of the shear pressure $\phi$. Regardless of the sign of $\phi$, almost trivial inequalities of the EoS and the transport coefficients are obtained from the sufficient conditions:
@@ III.2 Constraint on the inverse Reynolds number
By employing conformal EoS and transport coefficients from AdS/CFT, the causality conditions are written as inequalities among thermodynamic variables and the shear pressure: [^3] Although the pressure $p$ is just equal to $e/3$ in the conformal EoS, we still keep $p$ in these inequalities since we compare these results with the one from lattice EoS in Sec. III.4.  [^4] Our result is consistent with the one obtained in Ref. Bemfica et al. (2021): $-4/5<\Lambda_{a}/(e+p)\leq 3/5$ with a different choice of relaxation time, $\tau_{\pi}T=5\eta/s$.
@@
For the necessary conditions, Eqs. (18) and (19) provide the constraint (31) for $\phi<0$ and $\phi>0$, respectively. On the other hand, the constraint (32) comes from Eqs. (III.1) and (III.1). See also Appendix B to understand how each necessary condition is obtained from a discussion of characteristic velocity.
Figure 2: Examples of the solutions of hydrodynamic equations (9) and (10) for several initial conditions and acausal (red) and causal (blue) regions from the necessary (31) and the sufficient (32) conditions, respectively. Note that, in between one cannot determine whether a solution is causal or not in between only from these conditions.
To see how the solutions for hydrodynamic equations of one-dimensional expanding fluids obey or violate the causality, the regions from these conditions (31) and (32) together with ten examples of solutions are shown in Fig. 2. When a solution of Eqs. (9) and (10) goes through the acausal region, it does not obey at least one necessary condition and, consequently, violates the relativistic causality. On the other hand, when a solution keeps its trajectory in the causal region during the time evolution, it obeys the all sufficient conditions (needless to say, the all necessary conditions, too) and, consequently, the causality of the dynamics is guaranteed. With only these necessary and sufficient conditions of nonlinear causality, it is not yet determined in between whether it is causal or not: All necessary conditions are satisfied, but some of the sufficient conditions are not satisfied. In any case, the dynamics of fluids tends to be acausal when the inverse Reynolds number is large and the system is far away from local equilibrium.